HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

ANALYSIS OF THE HOT BANDS OF METHYLAMINE: 3ν_{15}←ν_{15}, 4ν_{15}←ν_{15} AND ν_{9}←ν_{15}

\underline{I. GULACZYK} , M. KREGLEWSKI, Faculty of Chemistry, Adam Mickiewicz University in Poznan, Poland; I. KLEINER, Universite Paris Cite and Univ Paris

Est Creteil, CNRS, LISA, 75013, Paris, France

Methylamine is a molecule performing two large amplitude motions: CH_{3} internal rotation and NH_{2} inversion. These two large amplitude motions are strongly coupled and give rise to a rotation-inversion-torsion structure in the vibrational states. The rovibrational spectrum of the methylamine molecule has been extensively studied both experimentally and theoretically. The analyses of infrared bands such as inversion (ν_{9}) or CN stretching (ν_{8}) show significant perturbations from highly excited torsional states. In order to untangle the interactions in the 700 - 1200 cm^{−1} region of the methylamine spectrum, it is crucial to assign the dark states, which are the third and fourth excited torsional states (3ν_{15} and 4ν_{15}). These states can be experimentally observed as hot bands from the first excited torsional bands in the range 360 - 720 cm^{−1}. In the same range a hot band to the excited inversion state was discovered and since the intensities of three hot bands are similar, they must be analyzed jointly. The spectra were recorded with a resolution of 0.00125 cm^{−1} using Bruker IFS-120HR spectrometer at the University of Oulu. The accurate energy levels of the first excited torsional state, ν_{15},^{1}^{2} were used as reference values for lower state combination differences in the assignments of the third and fourth torsional hot bands, 3ν_{15}←ν_{15}, 4ν_{15}←ν_{15}, as well as in the hot inversion-torsional band, ν_{9}←ν_{15}. On the basis of the predicted energy levels for the third and fourth excited torsional states, many transitions of the hot torsional bands: 3ν_{15}←ν_{15}, 4ν_{15}←ν_{15}, not assigned previously, have been identified. Many transitions belonging to the inversion-torsion hot band, ν_{9}←ν_{15}, have been assigned too. All the assignments were confirmed by the Lower State Combination Differences. Only experimental data of inversion-torsion hot band, ν_{9}←ν_{15}, could be fitted to a single state model based on the group theoretical formalism of Hougen and Ohashi^{3}. The transitions belonging to the hot torsional bands (3ν_{15}←ν_{15}, 4ν_{15}←ν_{15}) require another approach like a ”hybrid” Hamiltonian model^{4}^{5}, thus the analysis is still in progress.

1 I. Gulaczyk, M. Kreglewski, V-M. Horneman, J. Mol. Spectrosc. 342 (2017) 25-30.

2 I. Gulaczyk, M. Kreglewski, JQSRT 252 (2020) 107097.

3 N. Ohashi and J.T. Hougen, J. Mol. Spectrosc. 121, (1987) 474.

4 I. Kleiner, J.T. Hougen, J. Phys. Chem. A 119 (2015) 10664.

5 I. Kleiner, J.T. Hougen, J. Mol. Spectrosc. 368 (2020) 111255.

p-number: p038  Submitted on Tue, 06 Jun 2023 11:22:47 +0200